# Dynamic Conditional SKEPTIC

### v1

@@ 7.1 Model Fit and Diagnostic Checks
Using the full set of observations, we fit the DCC, DCC-NL, and DCS models to the S&P100 and S&P500 datasets. Model selection is evaluated according to the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC). For all the models, after having estimated the univariate volatility with a GARCH(1,1), we estimate the parameters $\phi=(\alpha,\beta)$. We remark that the estimate of the matrix ${Q}_{t}$ may not be positive semidefinite. In that case, we project it onto the nearest correlation matrix using the method of higham2002computing.
In the residual analysis, we evaluate the adequacy of model fitting to the data. A well-specified econometric model should effectively filter the observed data into a white noise series, uncorrelated over time. To assess the absence of correlation, we rely on the Portmanteau test (lutkepohl2005new), which examines the null hypothesis that residuals exhibit no autocorrelation up to lag $h$. The test statistic is defined as follows:
@@ 7.2 Portfolio Construction
    DCC-NL: the portfolio in (21), where $\widehat{H}_{t}$ is estimated from the unconditional correlation matrix processed with nonlinear shrinkage, this procedure follows engle2019large.
@@
    $1/p$: the equal-weighted portfolio, a common benchmark advocated by demiguel2009optimal.
For the estimation of the inverse correlation matrix, we start from a rank-based SKEPTIC estimator constructed from Kendall’s tau and Spearman’s rho statistics. Rank-based dependence measures are substantially more robust than Pearson correlation under heavy tails, skewness, and other departures from Gaussianity, and they provide superior support recovery for the precision matrix in high-dimensional settings liu2012high. To obtain a stable and well-conditioned estimate of the inverse correlation matrix, we impose sparsity through the QUIC algorithm of hsieh2014quic, selecting the regularization parameter $\lambda$ via the StARS stability criterion of liu2010stability [^4] The tuning parameter is set to $\lambda=0.596$ for the S&P100, and $\lambda=0.547$ for the S&P500.. This approach ensures that only the most reliable conditional dependencies are retained and avoids applying shrinkage to a Pearson-based correlation matrix, which is known to be less reliable in non-Gaussian environments.
@@ 7.3 Portfolio Backtesting
We set the confidence level at $\alpha=5\%$. To assess whether the model and estimation procedure produce reliable risk measure estimates, we conduct three backtesting procedures: the Unconditional Coverage (UC) test (kupiec1995techniques), the Conditional Coverage (CC) test (christoffersen2000relevant), and the Dynamic Quantile (DQ) test (engle2004caviar). These tests evaluate whether the empirical violation probability aligns with the expected violation probability implied by the confidence level, thereby capturing the realized risk. All three tests are based on the null hypothesis that the expected number of violations matches the observed number.
@@
Figure 1: Backtesting results for SP100 and SP500 across the DCC, DCC-NL, DCS-Rho, DCS-Tau and EW specifications. Return observations are plotted against their corresponding VaR and ES estimates, with violations highlighted in green.
As illustrated in Figure 1, the DCS-Tau and DCS-Rho models exhibit a more flexible and responsive behaviour than both the DCC-based models and the EW benchmark, enabling them to capture shocks in portfolio returns promptly. This flexibility translates into a superior ability to model and predict systemic risk—measured through VaR and Expected Shortfall—with the DCS-Tau model delivering the best overall performance.
@@ 8 Conclusion
We investigate the theoretical properties of the DCS model, including its stationarity and mixing properties, highlighting its robustness and flexibility in dynamically modeling the correlation matrix. The proposed model provides a more adaptable approach to capture complex market dynamics by addressing the limitations of the normality assumption in asset returns. A simulation study based on Montecarlo simulations shows the DCS in estimating the process parameters. The empirical analysis demonstrates that the DCS model achieves portfolio Sharpe ratios comparable to benchmark models while having lower turnover, thereby reducing portfolio costs. In the risk management application, the proposed estimators demonstrate optimal performance with a portfolio that, besides providing optimal average return and low rebalancing costs, ensures robust market risk tracking.
@@
where $F_{0,k}$ is the joint distribution function of $X_{0}$ and $X_{k}$, and $\|\cdot\|_{\mathrm{TV}}$ is total variation, in the Vitali sense. We revisit the result of beare2010copulas. From Sklar’s theorem, and liu2009nonparanormal we have
@@
where the eigenvalues $\left\{\lambda_{i}\right\}$ form a nonincreasing square-summable sequence of non-negative real numbers and the eigenfunctions $\left\{\phi_{i}\right\}$ form a complete orthonormal sequence in $L_{2}[0,1]$. We know from beare2010copulas that:
@@
We exploit the Lipschitz continuity of the estimators and align our proof to the results of han2018exponential. We recall that let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. Denote the Gaussian cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ by $\Phi_{i}$ and $\Phi_{j}$. The population version of Spearman’s rho is: $R^{\rho}]_{ij}=\operatorname{Corr}\left(\Phi_{i}\left(X_{i}\right),\Phi_{j}\left(X_{j}\right)\right)$, and the population version of Kendall’s tau is:
@@ D.3 Sharpe ratio Test
We perform the Sharpe ratio test in the ”PeerPerformance” R package described in ardia2018peer using the block bootstrap approach. Table 10 presents the results of the Sharpe ratio test.
@@
Table 10: Summary of the ardia2018peer test for Sharpe ratio. The table reports differences in the annualized Sharpe ratio values. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.

### v2

@@ 7.1 Model Fit and Diagnostic Checks
Using the full set of observations, we fit the DCC, DCC-NL, and DCS models to the S&P100 and S&P500 datasets. Model selection is evaluated according to the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC). For all the models, after having estimated the univariate volatility with a GARCH(1,1), we estimate the parameters $\phi=(\alpha,\beta)$. We remark that the estimate of the matrix ${Q}_{t}$ may not be positive semidefinite. In that case, we project it onto the nearest correlation matrix using the method of Higham [18].
In the residual analysis, we evaluate the adequacy of model fitting to the data. A well-specified econometric model should effectively filter the observed data into a white noise series, uncorrelated over time. To assess the absence of correlation, we rely on the Portmanteau test (27), which examines the null hypothesis that residuals exhibit no autocorrelation up to lag $h$. The test statistic is defined as follows:
@@ 7.2 Portfolio Construction
    DCC-NL: the portfolio in (20), where $\widehat{H}_{t}$ is estimated from the unconditional correlation matrix processed with nonlinear shrinkage, this procedure follows [10].
@@
    $1/p$: the equal-weighted portfolio, a common benchmark advocated by [9].
For the estimation of the inverse correlation matrix, we start from a rank-based SKEPTIC estimator constructed from Kendall’s tau and Spearman’s rho statistics. Rank-based dependence measures are substantially more robust than Pearson correlation under heavy tails, skewness, and other departures from Gaussianity, and they provide superior support recovery for the precision matrix in high-dimensional settings [24]. To obtain a stable and well-conditioned estimate of the inverse correlation matrix, we impose sparsity through the QUIC algorithm of Hsieh et al. [19], selecting the regularization parameter $\lambda$ via the StARS stability criterion of Liu et al. [26] [^4] The tuning parameter is set to $\lambda=0.596$ for the S&P100, and $\lambda=0.547$ for the S&P500.. This approach ensures that only the most reliable conditional dependencies are retained and avoids applying shrinkage to a Pearson-based correlation matrix, which is known to be less reliable in non-Gaussian environments.
@@ 7.3 Portfolio Backtesting
We set the confidence level at $\alpha=5\%$. To assess whether the model and estimation procedure produce reliable risk measure estimates, we conduct three backtesting procedures: the Unconditional Coverage (UC) test (Kupiec [21]), the Conditional Coverage (CC) test (Christoffersen and Diebold [5]), and the Dynamic Quantile (DQ) test (Engle and Manganelli [11]). These tests evaluate whether the empirical violation probability aligns with the expected violation probability implied by the confidence level, thereby capturing the realized risk. All three tests are based on the null hypothesis that the expected number of violations matches the observed number.
@@
Figure 1: Backtesting results for SP100 and SP500 across the DCC, DCC-NL, DCS (rho), DCS (tau) and EW specifications. Return observations are plotted against their corresponding VaR and ES estimates, with violations highlighted in green.
As illustrated in Figure 1, the DCS (tau) and DCS (rho) models exhibit a more flexible and responsive behaviour than both the DCC-based models and the EW benchmark, enabling them to capture shocks in portfolio returns promptly. This flexibility translates into a superior ability to model and predict systemic risk, measured through VaR and Expected Shortfall, with the DCS (tau) model delivering the best overall performance.
@@ 8 Conclusion
We investigate the theoretical properties of the DCS model, including its stationarity and mixing properties, highlighting its robustness and flexibility in dynamically modeling the correlation matrix. The proposed model provides a more adaptable approach to capture complex market dynamics by addressing the limitations of the normality assumption in asset returns. A Monte Carlo simulation study demonstrates the effectiveness of the DCS approach in estimating the process parameters. The empirical analysis demonstrates that the DCS model achieves portfolio Sharpe ratios comparable to benchmark models while having lower turnover, thereby reducing portfolio costs. In the risk management application, the proposed estimators demonstrate optimal performance with a portfolio that, besides providing optimal average return and low rebalancing costs, ensures robust market risk tracking.
@@
where $F_{0,k}$ is the joint distribution function of $X_{0}$ and $X_{k}$, and $\|\cdot\|_{\mathrm{TV}}$ is total variation, in the Vitali sense. We revisit the result of Beare [3]. From Sklar’s theorem, and Liu et al. [25] we have
@@
where the eigenvalues $\left\{\lambda_{i}\right\}$ form a nonincreasing square-summable sequence of non-negative real numbers and the eigenfunctions $\left\{\phi_{i}\right\}$ form a complete orthonormal sequence in $L_{2}[0,1]$. We know from Beare [3] that:
@@
We exploit the Lipschitz continuity of the estimators and align our proof to the results of Han [17]. We recall that let $\widetilde{X}_{i}$ and $\widetilde{X}_{j}$ be independent copies of the random variables $X_{i}$ and $X_{j}$, respectively. Denote the Gaussian cumulative distribution functions (CDFs) of $X_{i}$ and $X_{j}$ by $\Phi_{i}$ and $\Phi_{j}$. The population version of Spearman’s rho is: $R^{\rho}]_{ij}=\operatorname{Corr}\left(\Phi_{i}\left(X_{i}\right),\Phi_{j}\left(X_{j}\right)\right)$, and the population version of Kendall’s tau is:
@@ D.3 Sharpe ratio Test
We perform the Sharpe ratio test using the ”PeerPerformance” R package described in Ardia and Boudt [1], with a block bootstrap approach. Table 10 presents the results of the Sharpe ratio test.
@@
Table 10: Sharpe ratio test following Ardia and Boudt [1]. The table reports $\Delta$ Sharpe ratios (annualized), i.e., the difference between the two portfolios’ Sharpe ratios in each comparison. Note: $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.